# Exceptional Tannaka groups only arise from cubic threefolds

### v2

@@ 5.1. Statement
  • If $n=3$, then $X$ is either a cone over a nonnormal cubic surface or a twisted plane as in example 5.2.
@@ 5.2. Reduction to surfaces
Let $X\subset\mathbb{P}^{n+1}$ be a nonnormal integral cubic hypersurface with $n\geqslant 1$, then $X^{\mathrm{sing}}$ is a linear subspace of dimension $n-1$.
@@
We argue by induction on $n\geqslant 1$. For $n=1$ the only nonnormal integral cubic plane curves are the nodal and the cuspidal plane cubics, which both have only one singular point. Suppose $n\geqslant 2$. By Serre’s criterion of normality the singular locus of $X$ has dimension $n-1$. Let $Y\subset X^{\mathrm{sing}}$ be an irreducible component of dimension $n-1$ and suppose that $Y$ is not a linear subspace. Pick a point $y\in Y$ such that the line through $y$ and a general point $y^{\prime}\in Y\smallsetminus\{y\}$ is not contained in $Y$. As $\deg X=3$ and $Y\subset X^{\mathrm{sing}}$, every such line is contained in $X$. An open subset of $X$ is thus covered by lines going through $y$. It follows that $X$ is a cone with vertex $y$ over an integral cubic hypersurface $Z\subset\mathbb{P}^{n}$, thus $Y^{\mathrm{sing}}$ is a cone with vertex $y$ over $Z^{\mathrm{sing}}$. In particular we have $\dim Z^{\mathrm{sing}}=n-2$, hence the cubic hypersurface $Z$ is not normal. By the induction hypothesis, $Z^{\mathrm{sing}}$ is a linear subspace of dimension $n-2$, thus $Y^{\mathrm{sing}}$ is a linear subspace of dimension $n-1$. ∎
@@
    If $n=3$, then $X$ is either a cone over a nonnormal cubic surface or the twisted plane from example 5.2.
@@
By lemma 5.4 the singular locus is linear subspace of dimension $n-1$, which we may assume to have equations $x_{0}=x_{1}=0$. In such coordinates, the hypersurface $X$ is defined by an equation of the form
@@
for some linear forms $f_{1},f_{2},f_{3}$. Let $m+2$ be the dimension of the linear span of the linear forms $x_{0}$$x_{1}$$f_{1}$$f_{2}$ and $f_{3}$. If $n>m$, then $X$ is the cone over a cubic hypersurface in $\mathbb{P}^{m+1}$ which is necessarily integral and nonnormal. Since $m\leqslant 3\leqslant n$ the equality $n=m$ holds only for $m=3$ in which case we may choose $f_{1},f_{2},f_{3}$ as coordinate functions and obtain that $X$ is the twisted plane in example 5.2. ∎
With the above lemma, the proof of proposition 5.3 will be reduced at the end of the next section to the study of nonnormal cubic surfaces given below.
@@ 5.3. Nonnormal cubic surfaces
In order to conclude the proof of proposition 5.3 we need to describe nonnormal cubic surfaces which are not cones over nonnormal plane cubic curves. To do this, for a point $v\in\mathbb{P}(V^{\vee})\smallsetminus S$ let $W_{v}\subset V$ be the subspace of linear forms vanishing at $v$. Let
@@
be the corresponding projection. The geometry of the projection $p_{v}\colon S\to p_{v}(S)$ depends on the image of the point $v$ via the projection with center $E$ that we denote by
@@
$C\;:=q(S)\quad\subset\;\mathbb{P}(V^{\prime\vee})$
@@
The situation is resumed in fig. 1.
[…]
$B$$X_{v}^{\mathrm{sing}}$$p_{v}$
Figure 1. The two possibilities for $p_{v}\colon S\to X$ in proposition 5.6: The first line shows the case $q(v)\in C$, the second the case $q(v)\not\in C$.
@@
The nonnormality of $X_{v}$ follows from the description of $p_{v}\colon S\to X_{v}$. Indeed, if $X_{v}$ were to be normal, then $p_{v}$ would be an isomorphism because $p_{v}$ is finite birational and $S$ is smooth. But we will show that $p_{v}$ is not an isomorphism, thus the surface $X_{v}$ cannot be normal. The line bundle $\mathcal{O}_{S}(2)(-E)$ has self-interesction $3$ thus $S\subset\mathbb{P}(V^{\vee})$ has degree $3$ and the same is true for $X_{v}\subset\mathbb{P}(W_{v}^{\vee})$. By lemma 5.4 the singular locus $X_{v}^{\mathrm{sing}}$ is a line.
@@
We are now ready to conclude the proof. Let $P$ be the plane containing $v$ and the strict transform $B$ of a line in $\mathbb{P}^{2}$ not passing through $o$. Then $B\subset P$ is a smooth conic and any line through $v$ meets $B$ in two points (taking in account multiplicities). This shows that $B\to X_{v}$ is of degree $2$ onto its image. To conclude it suffices to show that $p_{v}\colon S\smallsetminus B\to X_{v}$ takes values in $X_{v}\smallsetminus p_{v}(B)$ and is injective. To do this, let $x,y\in S$ be distinct points such that $p_{v}(x)=p_{v}(y)$ and let $D$ be the line joining $x$$y$ and $v$. Note that $x$ and $y$ cannot lie both in the exceptional divisor, otherwise we would have $D=E$ hence $v\in S$. It follows that $x$ and $y$ have distinct images under the blow-up map $\pi\colon S\to\mathbb{P}^{2}$. Let $L\subset\mathbb{P}^{2}$ be the line joining $\pi(x)$ and $\pi(y)$, and let $Q\subset S$ be its strict transform. If the blow-up point $o$ were to belong to $L$, then $Q$ would be mapped to a line in $\mathbb{P}(V^{\vee})$ containing $x$ and $y$, implying $D=Q$ and $v\in S$ again. Therefore $L$ does not contain $o$, so $Q$ is embedded as a smooth conic in $\mathbb{P}(V^{\vee})$. To conclude, notice that the plane $P^{\prime}$ determined by $Q$ must be $P$: Otherwise, as argued above, $P$ and $P^{\prime}$ would meet in a single point which belongs to $S$; but $v\in P\cap P^{\prime}$ so that would imply $v\in S$ once more. Finally, we have $P=P^{\prime}$ hence $Q=B$. ∎
Proof of proposition 5.3.
We choose homogeneous coordinates $x_{0},x_{1},x_{2}$ on $\mathbb{P}^{2}$ and identify $S$ with the blow-up of $\mathbb{P}^{2}$ in $o=[1:0:0]$. Then a basis of $V$ is given by the monomials
$u_{ij}=x_{i}x_{j}\qquad\textup{with}\quad 0\leqslant i\leqslant j\leqslant 2\quad\textup{and}\quad(i,j)\neq(0,0),$
and $S$ is cut out by the vanishing of the polynomials
$h_{1}$ $=u_{02}u_{11}-u_{01}u_{12},$ $h_{2}$ $=u_{01}u_{22}-u_{02}u_{12},$ $h_{3}$ $=u_{11}u_{22}-u_{12}^{2}.$
With this notation the vector space $V^{\prime}$ is spanned by $u_{11}$$u_{12}$ and $u_{22}$ and the conic $C$ has equation $h_{3}=0$. If $v\in\mathbb{P}(V)$ is defined by the vanishing of $u_{01}$$u_{12}$$u_{22}$ and $u_{02}+u_{11}$, then $q(v)\in C$ and
@@
The discussion preceding theorem 9.2.1 in [Dol12] shows that up to a linear change of variables, any nonnormal integral cubic surface $X\subset\mathbb{P}^{3}$ which is not a cone over a nonnormal plane cubic curve has equation (5.1) or (5.2). Together with lemma 5.5 and proposition 5.6 this concludes the proof. ∎
@@ 6. On Fano varieties of lines
We now prove some facts about Fano varieties of lines that will be used for B. Let $k$ be an algebraically closed field. For a finite-dimensional $k$-vector space $V$, the Fano variety of lines on a subvariety $X\subset\mathbb{P}(V)$ is the $k$-scheme $F_{X}$ whose points with values in a $k$-scheme $S$ are rank $2$ vector bundles $\mathcal{V}\subset V\otimes_{k}\mathcal{O}_{S}$ on $S$ with locally free quotient such that $\mathbb{P}(\mathcal{V})\subset\mathbb{P}(V)\times S$ is contained in $X\times S$. Equivalently, $F_{X}$ is the Hilbert scheme parametrizing subvarieties of $X$ with Hilbert polynomial $P(z)=z+1$ with respect to line bundle $\mathcal{O}(1)$.
@@ 6.1. Rational normal scrolls
where $V(a_{i})=\textup{H}^{0}(\mathbb{P}^{1},\mathcal{O}_{\mathbb{P}^{1}}(a_{i}))$. The morphism $f_{a}$ is birational and embeds each fiber of $\pi$ linearly. Moreover, $f_{a}$ is a closed immersion if and only if $a_{i}>0$ for all $i=1,\dots,r$. The image of $f_{a}$ is called the rational normal scroll of type $a$ and will be denoted $S(a)$. The rational normal scroll $S(a)$ is smooth in the following cases [EH87, p. 6]:
@@
  • $a=(3)$. In this case $S(3)\subset\mathbb{P}^{3}$ is the twisted cubic, i.e. the triple Veronese embedding of the projective line. Hence
    $F(3)=\varnothing.$
@@
    $F(1,2)_{v}$ =$=\{\textup{fibers of the projection\leavevmode\nobreak\ $S\to E$}\},$
    $F(1,2)_{h}$ =$=\{E\}.$
@@ 6.2. The Fano variety of a cone
is surjective and smooth. Moreover, given a plane $E\subset W$, the closure of the fiber of $\pi_{0}$ in the corresponding point $[E]\in\operatorname{Gr}_{2}(W)$ is $\operatorname{Gr}_{2}(\pi^{-1}(E))$.
@@
(1) The subset $\pi_{1}^{-1}(Y)\subset F_{X}$ parametrizes lines contained in the closure of fibers of $X\dashrightarrow Y$. By hypothesis any $y\in Y$ belongs to some line $L\subset Y$. Therefore lines in the closure of fiber of $X\dashrightarrow Y$ in $y$ are contained in $\operatorname{Gr}_{2}(\pi^{-1}(E))=\overline{\pi_{0}^{-1}([E])}$ where $L=\mathbb{P}(E)$ and $[E]\in\operatorname{Gr}_{2}(W)$ is the point defined by $E$.
@@
It will be useful to combine proposition 6.3 with the following remark: For each irreducible component $F^{\prime}\subset F_{Y}$ and $Y^{\prime}\subset Y$, we have
@@
Suppose $\dim W=3$ and that $Y$ is an integral cubic plane curve. Then the preceding statement says that the Fano variety of lines in the cone $X$ is irreducible of dimension $12(\dim V-4)+$ 1.
@@
Suppose $\dim W=4$ and that $Y\subset\mathbb{P}(W)$ is a nonnormal integral cubic surface that is not a cone over a cubic plane curve. Borrow notation from section 5.3. By proposition 5.3 the cubic surface $Y$ is obtained from the blow-up $S$ of $\mathbb{P}^{2}$ in a point,
@@
by projecting it from a point $v\in\mathbb{P}(V^{\vee})$ not in $S$. Recall from example 6.2 that
@@
It follows from proposition 5.6 that such a projection induces a morphism between Fano varieties of lines $f\colon F_{S}\to F_{Y}$. The morphism $f$ is a closed embedding once restricted to irreducible components. Moreover, following the case distinction in proposition 5.6 there are two possibilities:
@@
  • otherwise $F_{Y,{\mathrm{red}}}=f(F_{S,v})\sqcup f(F_{S,h})\sqcup\{Y^{\mathrm{sing}}\}$ has three irreducible components (recall by lemma 5.4 that the singular locus $Y^{\mathrm{sing}}$ of $Y$ is a line).
The surface $Y$ is covered by lines, thus by proposition 6.3, the morphism $\pi_{0}$ sets up a bijection between the irreducible components of the Fano variety of $X$ and that of $Y$. Continuing with the above distinction, we have the following cases:
@@ 6.3. The Fano variety of singular cubic hypersurfaces
If $n=3$ and $\dim F\leqslant 2$, then, for all integer $i\geqslant 0$,
@@
When $X$ has isolated singularities, this is equation (1.21.1) in [AK77], whose proof relies on corollary 1.4, lemma 1.14 and prop. 1.15 (i) in loc. cit.. However, with the notation therein, the proof of these facts only uses that $F$ is the zero subscheme of a regular section of the locally free sheaf $\operatorname{Sym}^{3}Q$. By theorem 3.3 (iv) in loc. cit., this latter fact is true more generally when $\dim F\leqslant 2$, hence the result follows. Note that the statement of theorem 3.3 (iv) gives a somewhat weaker statement: for each $x\in F$ at which $F$ has dimension $\leqslant 2$, the subvariety $F$ is the zero subscheme of a section $s_{x}$ of $\operatorname{Sym}^{3}Q$ which is regular at $x$. Anyway, it is clear from its proof that the section $s_{x}$ can be taken to be the same for all points $x\in F$. ∎
The main goal of this section is to complement [AK77] by describing precisely the cases where $F$ is not of dimension $2$. When $X$ is a cone over a cubic surface with finitely many lines $F$ has dimension $2$; see example 6.5. On the other hand, in examples 6.6, LABEL: and 6.4 we saw that the Fano variety of a cone over a nonnormal cubic surface or over an integral cubic plane curve has dimension $3$. The twisted plane in example 5.2 also furnishes such an example:
@@
Let $X\subset\mathbb{P}^{4}$ be the twisted plane from example 5.2, given by
@@
  • one $3$-dimensional component is given by the lines in the $1$-dimensional family of planes in $X$ with equations, for $[t_{0}:t_{1}]\in\mathbb{P}^{1}$,
@@
  • one $2$-dimensional component obtained as the closure of the image of the morphism $\mathbb{P}^{2}\smallsetminus\{[0:1:0],[0:0:1]\}\to\operatorname{Gr}_{2}(5)$ associating to $[t_{0}:t_{1}:t_{2}]$ the line joining the points
    $[t_{0}^{2}:t_{1}t_{2}:-t_{0}t_{2}:0:t_{0}t_{1}],\quad[0:t_{1}^{2}:-2t_{0}t_{1}:t_{0}^{2}:0]\quad\in\;\mathbb{P}^{4}.$
@@
  • the twisted plane in example 6.8, or
@@
This is the combination of the following two lemmas with the classification of nonnormal cubic hypersurfaces in proposition 5.3.
@@
Let $x\in X$ and $F_{X,x}\subset F_{X}$ the subvariety of lines passing through $x$. If $X$ is not a cone, then
$\dim F_{X,x}\leqslant n-2.$
@@
Suppose that $X$ is not a cone. By [AK77, th. 4.2] the variety $F$ is smooth of dimension $42n-$ 4 at the point corresponding to any line not meeting the singular locus. Therefore, it remains to treat the case of lines going through singular points. Since $\dim X^{\mathrm{sing}}\leqslant n-2$ by hypothesis, lemma 6.10 implies that $F$ at a line going through a singular point has dimension $\leqslant 2(n-2)$. This concludes the proof. ∎
Proof of proposition 6.9.

### v3

@@ 5.1. Statement
  • If $n=3$, then $X$ is either a cone over a nonnormal cubic surface or a twisted plane as in example˜5.2.
@@ 5.2. Reduction to surfaces
Let $X\subset\mathbb{P}^{n+1}$ be a nonnormal integral cubic hypersurface with $n\geqslant 1$, then $X^{\mathrm{sing}}$ is a linear subspace of dimension n1n-1.
@@
We argue by induction on $n\geqslant 1$. For $n=1$ the only nonnormal integral cubic plane curves are the nodal and the cuspidal plane cubics, which both have only one singular point. Suppose $n\geqslant 2$. By Serre’s criterion of normality the singular locus of $X$ has dimension n1n-1. Let $Y\subset X^{\mathrm{sing}}$ be an irreducible component of dimension n1n-1 and suppose that $Y$ is not a linear subspace. Pick a point $y\in Y$ such that the line through $y$ and a general point $y^{\prime}\in Y\smallsetminus\{y\}$ is not contained in $Y$. As $\deg X=3$ and $Y\subset X^{\mathrm{sing}}$, every such line is contained in $X$. An open subset of $X$ is thus covered by lines going through $y$. It follows that $X$ is a cone with vertex $y$ over an integral cubic hypersurface $Z\subset\mathbb{P}^{n}$, thus $Y^{\mathrm{sing}}$ is a cone with vertex $y$ over $Z^{\mathrm{sing}}$. In particular we have $\dim Z^{\mathrm{sing}}=n-2$, hence the cubic hypersurface $Z$ is not normal. By the induction hypothesis, $Z^{\mathrm{sing}}$ is a linear subspace of dimension n2n-2, thus $Y^{\mathrm{sing}}$ is a linear subspace of dimension n1n-1. ∎
@@
    If $n=3$, then $X$ is either a cone over a nonnormal cubic surface or the twisted plane from example˜5.2.
@@
By lemma˜5.4, the singular locus is a linear subspace of dimension n1n-1, which we may assume to have equations $x_{0}=x_{1}=0$. In such coordinates, the hypersurface $X$ is defined by an equation of the form
@@
for some linear forms $f_{1},f_{2},f_{3}$. Let $m+2$ be the dimension of the linear span of the linear forms $x_{0}$, $x_{1}$, $f_{1}$, $f_{2}$ and $f_{3}$. If $n>m$, then $X$ is the cone over a cubic hypersurface in $\mathbb{P}^{m+1}$ which is necessarily integral and nonnormal. Since $m\leqslant 3\leqslant n$ the equality $n=m$ holds only for $m=3$ in which case we may choose $f_{1},f_{2},f_{3}$ as coordinate functions and obtain that $X$ is the twisted plane in example˜5.2. ∎
With the above lemma, the proof of proposition˜5.3 will be reduced at the end of the next section to the study of nonnormal cubic surfaces given below.
@@ 5.3. Nonnormal cubic surfaces
In order to conclude the proof of proposition˜5.3 we need to describe nonnormal cubic surfaces which are not cones over nonnormal plane cubic curves. To do this, for a point $v\in\mathbb{P}(V^{\vee})\smallsetminus S$ let $W_{v}\subset V$ be the subspace of linear forms vanishing at $v$. Let
@@
be the corresponding projection. The geometry of the projection $p_{v}\colon S\to p_{v}(S)$ depends on the image of $v$ via the projection with center $E$ that we denote by
@@
$C\;:=q(S)\;\subset\;\mathbb{P}(V^{\prime\vee})$
@@
The nonnormality of $X_{v}$ follows from the description of $p_{v}\colon S\to X_{v}$. Indeed, if $X_{v}$ were to be normal, then $p_{v}$ would be an isomorphism because $p_{v}$ is finite birational and $S$ is smooth. But we will show that $p_{v}$ is not an isomorphism, thus the surface $X_{v}$ cannot be normal. The line bundle $\mathcal{O}_{S}(2)(-E)$ has self-intersection $3$; thus $S\subset\mathbb{P}(V^{\vee})$ has degree $3$ and the same is true for $X_{v}\subset\mathbb{P}(W_{v}^{\vee})$. By lemma˜5.4 the singular locus $X_{v}^{\mathrm{sing}}$ is a line.
@@
We are now ready to conclude the proof. Let $P$ be the plane containing $v$ and the strict transform $B$ of a line in $\mathbb{P}^{2}$ not passing through $o$. Then $B\subset P$ is a smooth conic and any line through $v$ meets $B$ in two points (taking in account multiplicities). This shows that $B\to X_{v}$ is of degree $2$ onto its image. To conclude it suffices to show that $p_{v}\colon S\smallsetminus B\to X_{v}$ takes values in $X_{v}\smallsetminus p_{v}(B)$ and is injective. To do this, let $x,y\in S$ be distinct points such that $p_{v}(x)=p_{v}(y)$ and let $D$ be the line joining $x$, $y$ and $v$. Note that $x$ and $y$ cannot lie both in the exceptional divisor, otherwise we would have $D=E$ hence $v\in S$. It follows that $x$ and $y$ have distinct images under the blow-up map $\pi\colon S\to\mathbb{P}^{2}$. Let $L\subset\mathbb{P}^{2}$ be the line joining $\pi(x)$ and $\pi(y)$, and let $Q\subset S$ be its strict transform. If the blow-up point $o$ were to belong to $L$, then $Q$ would be mapped to a line in $\mathbb{P}(V^{\vee})$ containing $x$ and $y$, implying $D=Q$ and $v\in S$ again. Therefore $L$ does not contain $o$, so $Q$ is embedded as a smooth conic in $\mathbb{P}(V^{\vee})$. To conclude, notice that the plane $P^{\prime}$ determined by $Q$ must be $P$: Otherwise, as argued above, $P$ and $P^{\prime}$ would meet in a single point which belongs to $S$; but $v\in P\cap P^{\prime}$ so that would imply $v\in S$ once more. Finally, we have $P=P^{\prime}$ hence $Q=B$. ∎
Proof of proposition˜5.3.
We choose homogeneous coordinates $x_{0},x_{1},x_{2}$ on $\mathbb{P}^{2}$ and identify $S$ with the blow-up of $\mathbb{P}^{2}$ in $o=[1:0:0]$. Then a basis of $V$ is given by the monomials $u_{ij}=x_{i}x_{j}$ with $0\leqslant i\leqslant j\leqslant 2$ and $(i,j)\neq(0,0)$, and $S$ is cut out by the vanishing of the polynomials $h_{1}=u_{02}u_{11}-u_{01}u_{12}$, $h_{2}=u_{01}u_{22}-u_{02}u_{12}$ and $h_{3}=u_{11}u_{22}-u_{12}^{2}$. With this notation the vector space $V^{\prime}$ is spanned by $u_{11}$, $u_{12}$ and $u_{22}$ and the conic $C$ has equation $h_{3}=0$. If $v\in\mathbb{P}(V)$ is defined by the vanishing of $u_{01}$, $u_{12}$, $u_{22}$ and $u_{02}+u_{11}$, then $q(v)\in C$ and
@@
The discussion preceding theorem 9.2.1 in [Dol12] shows that up to a linear change of variables, any nonnormal integral cubic surface $X\subset\mathbb{P}^{3}$ which is not a cone over a nonnormal plane cubic curve has equation (5.1) or (5.2). Together with lemma˜5.5 and proposition˜5.6 this concludes the proof. ∎
@@ 6. On Fano varieties of lines
We now prove some facts about Fano varieties of lines that will be used for ˜B. Let $k$ be an algebraically closed field. For a finite-dimensional $k$-vector space $V$, the Fano variety of lines on a subvariety $X\subset\mathbb{P}(V)$ is the $k$-scheme $F_{X}$ whose points with values in a $k$-scheme $S$ are rank $2$ vector bundles $\mathcal{V}\subset V\otimes_{k}\mathcal{O}_{S}$ on $S$ with locally free quotient such that $\mathbb{P}(\mathcal{V})\subset\mathbb{P}(V)\times S$ is contained in $X\times S$. Equivalently, $F_{X}$ is the Hilbert scheme parametrizing subvarieties of $X$ with Hilbert polynomial $P(z)=z+1$ with respect to line bundle $\mathcal{O}(1)$.
@@ 6.1. Rational normal scrolls
where $V(a_{i})=\textup{H}^{0}(\mathbb{P}^{1},\mathcal{O}_{\mathbb{P}^{1}}(a_{i}))$. The morphism $f_{a}$ is birational and embeds each fiber $F$ of $p_{a}$ in such a way that $f_{a}^{\ast}\mathcal{O}(1)\simeq\mathcal{O}_{F}(1)$. Moreover, $f_{a}$ is a closed immersion if and only if $a_{i}>0$ for all $i=1,\dots,r$. The image of $f_{a}$ is called the rational normal scroll of type $a$ and will be denoted $S(a)$. The rational normal scroll $S(a)$ is smooth in the following cases [EH87, p. 6]:
@@
  • $a=(3)$. In this case $S(3)\subset\mathbb{P}^{3}$ is the twisted cubic, i.e. the triple Veronese embedding of the projective line. Hence $F(3)=\varnothing$.
@@
    $F(1,2)_{v}=\{\textup{fibers of the projection\penalty 10000\ $S\to E$}\},\qquad F(1,2)_{h}=\{E\}.$
@@ 6.2. The Fano variety of a cone
is surjective and smooth. Moreover, given a plane $E\subset W$, the closure of the fiber of $\pi_{0}$ in the corresponding point $E]\in\operatorname{Gr}_{2}(W)$ is $\operatorname{Gr}_{2}(\pi^{-1}(E))$.
@@
(1) The subset $\pi_{1}^{-1}(Y)\subset F_{X}$ parametrizes lines contained in the closure of fibers of $X\dashrightarrow Y$. By hypothesis any $y\in Y$ belongs to some line $L\subset Y$. Therefore lines in the closure of fiber of $X\dashrightarrow Y$ in $y$ are contained in $\operatorname{Gr}_{2}(\pi^{-1}(E))=\overline{\pi_{0}^{-1}([E])}$ where $L=\mathbb{P}(E)$ and $E]\in\operatorname{Gr}_{2}(W)$ is the point defined by $E$.
@@
It will be useful to combine proposition˜6.3 with the following remark: For each irreducible component $F^{\prime}\subset F_{Y}$ and $Y^{\prime}\subset Y$, we have
@@
Suppose $\dim W=3$ and that $Y$ is an integral cubic plane curve. Then the preceding statement says that the Fano variety of lines in the cone $X$ is irreducible of dimension $2(\dim V-4)+1$.
@@
Suppose $\dim W=4$ and that $Y\subset\mathbb{P}(W)$ is a nonnormal integral cubic surface that is not a cone over a cubic plane curve. Borrow notation from section˜5.3. By proposition˜5.3 the cubic surface $Y$ is obtained from the blow-up $S$ of $\mathbb{P}^{2}$ in a point,
@@
by projecting it from a point $v\in\mathbb{P}(V^{\vee})$ not in $S$. Recall from example˜6.2 that
@@
It follows from proposition˜5.6 that such a projection induces a morphism between Fano varieties of lines $f\colon F_{S}\to F_{Y}$. The morphism $f$ is a closed embedding once restricted to irreducible components. Moreover, following the case distinction in proposition˜5.6 there are two possibilities:
@@
  • otherwise $F_{Y,{\mathrm{red}}}=f(F_{S,v})\sqcup f(F_{S,h})\sqcup\{Y^{\mathrm{sing}}\}$ has three irreducible components (recall by lemma˜5.4 that the singular locus $Y^{\mathrm{sing}}$ of $Y$ is a line).
The surface $Y$ is covered by lines, thus by proposition˜6.3, the morphism $\pi_{0}$ sets up a bijection between the irreducible components of the Fano variety of $X$ and that of $Y$. Continuing with the above distinction, we have the following cases:
@@ 6.3. The Fano variety of singular cubic hypersurfaces
If $n=3$ and $\dim F\leqslant 2$, then $F$ is a locally complete intersection of pure dimension $2$, and for all $i\geqslant 0$,
@@
With the notation in [AK77], theorem 3.3 (iv) therein states that $F$ is the zero locus of a regular section of the locally free sheaf $\operatorname{Sym}^{3}Q$. Note that the statement of theorem 3.3 (iv) gives a somewhat weaker statement: for each $x\in F$ at which $F$ has dimension $\leqslant 2$, the subvariety $F$ is the zero subscheme of a section $s_{x}$ of $\operatorname{Sym}^{3}Q$ which is regular at $x$. Anyway, it is clear from its proof that the section $s_{x}$ can be taken to be the same for all points $x\in F$. When $X$ has isolated singularities, the identity in the statement is equation (1.21.1) in [AK77], whose proof relies on corollary 1.4, lemma 1.14 and prop. 1.15 (i) in loc. cit.. However, the proof of these facts only uses that $F$ is the zero locus of a regular section of $\operatorname{Sym}^{3}Q$, hence holds under our more general assumptions. ∎
The main goal of this section is to complement [AK77] by describing precisely the cases where $F$ is not of dimension $2$. When $X$ is a cone over a cubic surface with finitely many lines $F$ has dimension $2$; see example˜6.5. On the other hand, in examples˜6.6, LABEL: and 6.4 we saw that the Fano variety of a cone over a nonnormal cubic surface or over an integral cubic plane curve has dimension $3$. The twisted plane in example˜5.2 also furnishes such an example:
@@
Let $X\subset\mathbb{P}^{4}$ be the twisted plane from example˜5.2, given by
@@
  • one $3$-dimensional component is given by the lines in the $1$-dimensional family of planes in $X$ with equations, for $t_{0}:t_{1}]\in\mathbb{P}^{1}$,
@@
  • one $2$-dimensional component obtained as the closure of the image of the morphism $\mathbb{P}^{2}\smallsetminus\{[0:1:0],[0:0:1]\}\to\operatorname{Gr}_{2}(5)$ associating to $t_{0}:t_{1}:t_{2}]$ the line joining the points
    $t_{0}^{2}:t_{1}t_{2}:-t_{0}t_{2}:0:t_{0}t_{1}],\quad[0:t_{1}^{2}:-2t_{0}t_{1}:t_{0}^{2}:0]\quad\in\;\mathbb{P}^{4}.$
@@
  • the twisted plane in example˜6.8, or
@@
This is the combination of the following two lemmas with the classification of nonnormal cubic hypersurfaces in proposition˜5.3.
@@
Let $x\in X$ and $F_{X,x}\subset F_{X}$ the subvariety of lines passing through $x$. If $X$ is not a cone, then $\dim F_{X,x}\leqslant n-2$.
@@
Suppose that $X$ is not a cone. By [AK77, th. 4.2] the variety $F$ is smooth of dimension 2n42n-4 at the point corresponding to any line not meeting the singular locus. Therefore, it remains to treat the case of lines going through singular points. Since $\dim X^{\mathrm{sing}}\leqslant n-2$ by hypothesis, lemma˜6.10 implies that $F$ at a line going through a singular point has dimension $\leqslant 2(n-2)$. This concludes the proof. ∎
Proof of proposition˜6.9.